Water waves in a small tank are .06 m long. They pass a given point at a rate of 14.8 waves every three seconds. What is the speed of the water waves? What is the period of these water waves?
1
Expert's answer
2019-10-22T10:24:45-0400

The frequency:


"f=\frac{14.8}{3}Hz"

The speed of the water waves:


"v=\lambda f=0.06\frac{14.8}{3}=0.296\frac{m}{s}"

The period of these water waves:


"T=\frac{3}{14.8}=0.203\ s"